# Optical systems with translational invariance

### v2

@@
Optical systems with translational invariance
Dmitry Zhuridov Institute of Theoretical Physics, University of Wrocław, Plac Maxa Borna 9, 50204, Wrocław, Poland dmitry.zhuridov@uwr.edu.pl
journal: opticajournalarticletype: Research Article{abstract*}
The leading order monochromatic aberrations are investigated for the optical systems, which obey single-plane symmetry and translational invariance. These aberrations are classified from the symmetry principles for the wave aberration function. The Fermat’s Principle is applied for a system made of generic cylindrical surfaces and a complete set of the aberration coefficients required for calculation of the aberrations for any paraxial ray is obtained. To demonstrate the applied value of the analytical results obtained the criteria of compensation of the leading aberrations for cylindrical analog of the Cassegrain telescope were found.
@@ 2 Aberrations for cylindrical surfaces: generic picture
Number of aberrations for generic optical surfaces is high. Wave aberration $\Phi$ is a function of four parameters [12], e.g., the coordinates {$x$, $y$, $x^{\prime}$, $y^{\prime}$} of the points $P_{0}(x,y,0)$ and $P^{\prime}(x^{\prime},y^{\prime},z^{\prime})$, in which a paraxial ray crosses the object plane and the exit pupil plane, respectively. These two planes are orthogonal to the $Z$ axis (in the case of rotatoinally symmetric OS it is the symmetry axis, while in the case of discussed translationally invariant OS it can be any selected axis that is orthogonal to the object plane and to the vector of translation). Expansion of $\Phi$ into the four argument power series gives 35 terms in the fourth-order part $\Phi^{(4)}$ that determines the number of the third-order aberrations. (We notice that the four terms, which are proportional to $x^{3}$, $x^{2}y$, $xy^{2}$ and $y^{3}$, reduce the number of the third-order aberrations by four units, since they do not contribute to Eqs. (8) and (9) that are given below.) Number of these aberrations can be reduced for OS with definite symmetries, in particular, single-plane or double-plane symmetry. In the paraxial region a generic anamorphic OS (double-plane symmetric one) can be replaced by two independent rotationally symmetric OS, each associated with one symmetry plane [1]. This helps to derive 16 monochromatic primary aberration coefficients for anamorphic systems made from cylindrical surfaces [2].
Symmetry of cylindrical surface, namely, single-plane symmetry plus translational invariance, allows for a simplified generic consideration and explicit analytical description presented in the following. Consider a system of cylindrical surfaces, which is symmetric under reflection in the $YZ$ plane and translationally invariant along the $Y$ axis. The translational invariance leaves three independent variables, e.g., $x$, $x^{\prime}$ and the difference $y-y^{\prime}$, from which one can construct the six invariants of the second order:
@@
The discussed plane symmetry leaves the following four invariants:
@@
which completely determine the ray position. Since the variables enter only in the quadratic combinations of Eq. (2), it follows that $\Phi^{(4)}$ must be of the form
$\Phi^{(4)}$ $=$ $a_{1}A^{2}+\frac{a_{2}}{4}B^{2}+\frac{a_{3}}{4}C^{2}+\frac{a_{4}}{2}D^{2}$ (3)
$+$ $\frac{a_{5}}{2}AB+\frac{a_{6}}{2}AC+a_{7}AD+\frac{a_{8}}{2}BC+a_{9}BD+a_{10}CD,$
where $a_{i}$ are the constants, and the numerical factors $1/2$ and $1/4$ are introduced for the further convenience.
@@
where $O^{\prime}P_{1}^{*}=R^{\prime}$ is the radius of this sphere and $D_{1}$ is the distance between the exit pupil plane and the image plane. Following the consideration of Ref. [13], we take into account the fact that for the case of small aberrations in the paraxial region the directions of normal to the Gaussian reference sphere and normal to the real wave surface for the respective actual ray $P^{\prime}P_{1}$ are about the same. Then equation for the real wave surface can approximately be written as
$f(x^{\prime},y^{\prime},z^{\prime})\equiv(x^{\prime}-x_{1}^{*})^{2}+(y^{\prime}-y_{1}^{*})^{2}+(z^{\prime}-D_{1})^{2}-(s^{\prime}+\Phi)^{2}=0,$ (5)
where $O^{\prime}P_{1}=s^{\prime}$ is the distance from $O^{\prime}$ to the actual image point $P_{1}$, and $\Phi\ll R^{\prime}\approx s^{\prime}$. Equation of the normal to the real wave in the point with the coordinates $x^{\prime}$, $y^{\prime}$ and $z^{\prime}$ can be expressed as
@@
Substituting the coordinates of $P_{1}(x_{1},y_{1},D_{1})$ into Eq. (7), we get the following expressions for the components of transverse aberration
@@
$\frac{\delta_{1x}}{D_{1}}$ $=$ $(a_{2}B+a_{5}A+a_{8}C+2a_{9}D)x^{\prime}+(a_{4}D+a_{7}A+a_{9}B+a_{10}C)x,$ (10)
@@
$\delta_{1x}=(\tilde{a}_{2}x^{\prime 2}+\tilde{a}_{8}y^{\prime 2})x^{\prime}$ $+(3\tilde{a}_{9}x^{\prime 2}+\tilde{a}_{10}y^{\prime 2})x$ $+(\tilde{a}_{4}+\tilde{a}_{5})x^{2}x^{\prime}\quad+\tilde{a}_{7}x^{3},$ (12)
@@
where in the right-hand sides the first, second and third terms account for the three-component elliptic aberration ($e$), two-component modified coma ($c$) and astigmatism with the field curvature ($a$), respectively, while the last term in Eq. (12) represents the distortion ($d$). These aberrations are discussed one by one in the following. In total they contain 8 degrees of freedom since $\tilde{a}_{4}$ and $\tilde{a}_{5}$ parameters enter the same term. In case of $\tilde{a}_{2}=\tilde{a}_{3}=\tilde{a}_{8}$ and $\tilde{a}_{9}=\tilde{a}_{10}$ we end up with the standard set of five aberrations for the spherically symmetric OS.
@@ 2.1 Elliptic aberrations
The system of two elliptic aberrations depend on the three parameters $\tilde{a}_{2}$, $\tilde{a}_{3}$ and $\tilde{a}_{8}$. In the polar coordinates
@@
one can split Eqs. (16)-(17) into the two sets of parametric equations for the two ellipses:
@@
These two elliptic aberrations degenerate to the two spherical ones (with the opposite circle directions) in the case of $\tilde{a}_{2}=\tilde{a}_{3}$, which in turn degenerate to the one spherical aberration in the case of $\tilde{a}_{2}=\tilde{a}_{3}=\tilde{a}_{8}$.
@@ 2.2 Modified coma
In the same coordinates (20)-(21) the terms representing modified coma can be written as
@@ 2.3 Astigmatism and field curvature
This gives the same equation of ellipse as for the spherically symmetric systems:
@@ 2.4 Distortion
It does not depend on the coordinates $x^{\prime}$ and $y^{\prime}$ of the point, in which the ray crosses the exit pupil. Similarly to the case of spherically symmetric systems, the departure of the image point from the ideal image one due to the distortion is proportional to $x^{3}$.
@@ 3.1 Application of Fermat’s Principle
The homogeneous medium before the surface along the $Z$ axis has refraction index $n_{0}$ and the one after the surface has index $n_{1}$. We remark that Eq. (38) corresponds to the $XZ$ section of a generalized surface of rotation discussed in Ref. [14], where $\kappa$ is the conic constant and the $b$ term explicitly includes the type of aspheric term required for corrector plates. In particular, $\alpha=0$ for parabolic cylinder, $\alpha=R^{-3}$ for cylinder of revolution and $\alpha=(1+\kappa)R^{-3}$ with $\kappa<-1$ for hyperbolic cylinder.
@@
The Fig. 1 shows the cylindrical surface of Eq. (38), which generatrix coincides with the $Y$ axis and the symmetry axis of the aperture stop. The optical path length between the object point $P_{0}$ and the image point $P_{1}$ for the ray crossing the surface in $P^{\prime}$ can be written as
$n_{0}P_{0}P^{\prime}+n_{1}P^{\prime}P_{1}=$
$n_{0}\sqrt{(x^{\prime}-x_{0})^{2}+y^{\prime 2}+(z^{\prime}-z_{0})^{2}}+n_{1}\sqrt{(x^{\prime}-x_{1})^{2}+(y^{\prime}-y_{1})^{2}+(z^{\prime}-z_{1})^{2}},$ (39)
@@
Since normal to the surface in Eq. (38) in the points along the $Y$ axis is parallel to the $Z$ axis, the Snell’s law of refraction for the chief ray can be written as
@@
$P_{0}P^{\prime}=s_{0}\sqrt{1+\frac{2x^{\prime}\sin\theta_{0}}{s_{0}}+\frac{x^{\prime 2}}{s_{0}}\left(\frac{1}{s_{0}}+\frac{\cos\theta_{0}}{R}\right)+\frac{x^{\prime 4}}{4s_{0}^{2}}\left(\frac{1}{R^{2}}+\alpha s_{0}\cos\theta_{0}\right)+\frac{y^{\prime 2}}{s_{0}^{2}}}.$ (45)
@@
$P_{0}P^{\prime}$ $=$ $s_{0}+x^{\prime}\sin\theta_{0}+\frac{\Theta_{0}^{x}}{2}\left(\frac{x^{\prime 2}}{s_{0}}-\frac{x^{\prime 3}}{s_{0}^{2}}\sin\theta_{0}\right)+\frac{1}{2}\left(\frac{y^{\prime 2}}{s_{0}}-\frac{x^{\prime}y^{\prime 2}}{s_{0}^{2}}\sin\theta_{0}\right)$ (46)
-- $\frac{1}{8s_{0}^{3}}\left(x^{\prime 4}\Theta_{0}^{xx}-2x^{\prime 2}y^{\prime 2}\Theta_{0}^{xy}+y^{\prime 4}\right),$
@@
$\Theta_{0}^{x}$ $=$ $\cos\theta_{0}\left(\cos\theta_{0}+\frac{s_{0}}{R}\right),\qquad\quad\Theta_{0}^{xy}=2-3\cos^{2}\theta_{0}-\frac{s_{0}}{R}\cos\theta_{0},$ (47)
$\Theta_{0}^{xx}$ $=$ $5\sin^{4}\theta_{0}-6\sin^{2}\theta_{0}\Theta_{0}+\Theta_{0}^{2}-\frac{s_{0}^{2}}{R^{2}}-\alpha s_{0}^{3}\cos\theta_{0},\ \ \Theta_{0}=1+\frac{s_{0}}{R}\cos\theta_{0}.$ (48)
Using Eq. (44) and relation $\sin\eta\approx\tan\eta$ for small paraxial $\eta$, we express $y_{1}$ as
@@
where we ignored $O^{\prime}P^{\prime}$ with respect to $P_{0}P^{\prime}$. Using Eq. (41), we get
@@
$P^{\prime}P_{1}$ $=$ $s_{1}\sqrt{1-\frac{2x^{\prime}\sin\theta_{1}}{s_{1}}+\frac{x^{\prime 2}}{s_{1}}\left(\frac{1}{s_{1}}-\frac{\cos\theta_{1}}{R}\right)+\frac{x^{\prime 4}}{4s_{1}^{2}}\left(\frac{1}{R^{2}}-\alpha s_{1}\cos\theta_{1}\right)+\frac{y^{\prime 2}}{s_{1}^{2}}\beta^{2}}$ (52)
$=$ $s_{1}-x^{\prime}\sin\theta_{1}+\frac{\Theta_{1}^{x}}{2}\left(\frac{x^{\prime 2}}{s_{1}}+\frac{x^{\prime 3}}{s_{1}^{2}}\sin\theta_{1}\right)+\frac{\beta^{2}}{2}\left(\frac{y^{\prime 2}}{s_{1}}+\frac{x^{\prime}y^{\prime 2}}{s_{1}^{2}}\sin\theta_{1}\right)$
@@
$\Theta_{1}^{x}$ $=$ $\cos\theta_{1}\left(\cos\theta_{1}-\frac{s_{1}}{R}\right),\qquad\quad\Theta_{1}^{xy}=2-3\cos^{2}\theta_{1}+\frac{s_{1}}{R}\cos\theta_{1},$ (53)
$\Theta_{1}^{xx}$ $=$ $5\sin^{4}\theta_{1}-6\sin^{2}\theta_{1}\Theta_{1}+\Theta_{1}^{2}-\frac{s_{1}^{2}}{R^{2}}+\alpha s_{1}^{3}\cos\theta_{1},\ \ \Theta_{1}=1-\frac{s_{1}}{R}\cos\theta_{1}.$ (54)
Now the optical path length in Eq. (3.1) can be rewritten as
$(n_{0}s_{0}+n_{1}s_{1})+x^{\prime}(n_{0}\sin\theta_{0}-n_{1}\sin\theta_{1})+\frac{x^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}\Theta_{0}^{x}+\frac{n_{1}}{s_{1}}\Theta_{1}^{x}\right)$ (55)
-- $\frac{x^{\prime 3}}{2}\left(n_{0}\frac{\sin\theta_{0}}{s_{0}^{2}}\Theta_{0}^{x}-n_{1}\frac{\sin\theta_{1}}{s_{1}^{2}}\Theta_{1}^{x}\right)+\frac{y^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}+\frac{n_{1}\beta^{2}}{s_{1}}\right)-\frac{x^{\prime}y^{\prime 2}}{2}\left(\frac{n_{0}\sin\theta_{0}}{s_{0}^{2}}-\frac{n_{1}\beta^{2}\sin\theta_{1}}{s_{1}^{2}}\right)$
-- $\frac{x^{\prime 4}}{8}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xx}+\frac{n_{1}}{s_{1}^{3}}\Theta_{1}^{xx}\right)+\frac{x^{\prime 2}y^{\prime 2}}{4}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xy}+\frac{n_{1}\beta^{2}}{s_{1}^{3}}\Theta_{1}^{xy}\right)-\frac{y^{\prime 4}}{8}\left(\frac{n_{0}}{s_{0}^{3}}+\frac{n_{1}\beta^{4}}{s_{1}^{3}}\right).$
For the chief ray optical path length at the same order of precision we have
$n_{0}\sqrt{x_{0}^{2}+z_{0}^{2}}+n_{1}\sqrt{x_{1}^{2}+y_{1}^{2}+z_{1}^{2}}=n_{0}s_{0}+n_{1}s_{1}\sqrt{1+\frac{y_{1}^{2}}{s_{1}^{2}}}\approx(n_{0}s_{0}+n_{1}s_{1})+\frac{n_{1}y_{1}^{2}}{2s_{1}}-\frac{n_{1}y_{1}^{4}}{8s_{1}^{3}}$ (56)
$=(n_{0}s_{0}+n_{1}s_{1})+\frac{n_{1}y^{\prime 2}}{2s_{1}}\left(1+\beta\right)^{2}-\frac{n_{1}y^{\prime 4}}{8s_{1}^{3}}\left(1+\beta\right)^{4}.$
By subtracting Eq. (56) from Eq. (55) we get the wave aberration
$\Phi^{(4)}=x^{\prime}(n_{0}\sin\theta_{0}-n_{1}\sin\theta_{1})+\frac{x^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}\Theta_{0}^{x}+\frac{n_{1}}{s_{1}}\Theta_{1}^{x}\right)$
$-\frac{x^{\prime 3}}{2}\left(n_{0}\frac{\sin\theta_{0}}{s_{0}^{2}}\Theta_{0}^{x}-n_{1}\frac{\sin\theta_{1}}{s_{1}^{2}}\Theta_{1}^{x}\right)+\frac{y^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}-n_{1}\frac{1+2\beta}{s_{1}}\right)-\frac{x^{\prime}y^{\prime 2}}{2}\left(\frac{n_{0}\sin\theta_{0}}{s_{0}^{2}}-\frac{n_{1}\beta^{2}\sin\theta_{1}}{s_{1}^{2}}\right)$
$-\frac{x^{\prime 4}}{8}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xx}+\frac{n_{1}}{s_{1}^{3}}\Theta_{1}^{xx}\right)+\frac{x^{\prime 2}y^{\prime 2}}{4}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xy}+\frac{n_{1}\beta^{2}}{s_{1}^{3}}\Theta_{1}^{xy}\right)-\frac{y^{\prime 4}}{8}\left[\frac{n_{0}}{s_{0}^{3}}-\frac{n_{1}}{s_{1}^{3}}(1+4\beta+6\beta^{2}+4\beta^{3})\right].$
$=\alpha_{d}\,x^{\prime}+\alpha_{a1}\frac{x^{\prime 2}}{2}+\alpha_{a2}\frac{y^{\prime 2}}{2}+\alpha_{c1}\frac{x^{\prime 3}}{3}+\alpha_{c2}\,x^{\prime}y^{\prime 2}+\alpha_{e1}\frac{x^{\prime 4}}{4}+\alpha_{e2}\frac{y^{\prime 4}}{4}+\alpha_{e3}\frac{x^{\prime 2}y^{\prime 2}}{2},$ (57)
where
$\alpha_{e1}$ $=$ $-\frac{1}{2}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xx}+\frac{n_{1}}{s_{1}^{3}}\Theta_{1}^{xx}\right),\qquad\alpha_{e2}=-\frac{n_{0}}{2s_{0}^{3}}+\frac{n_{1}}{2s_{1}^{3}}(1+4\beta+6\beta^{2}+4\beta^{3}),$ (58)

### v3

@@
Abstract
The leading order monochromatic aberrations are investigated for the optical systems, which obey single-plane symmetry and translational invariance. These aberrations are classified from the symmetry principles for the wave aberration function. Fermat’s Principle is applied for a system made of generic cylindrical surfaces and a complete set of the aberration coefficients required for calculation of the aberrations for any paraxial ray is obtained. To demonstrate the applied value of the analytical results obtained the criteria of compensation of the leading aberrations for cylindrical analog of the Cassegrain telescope were found.
journal: opticajournal
@@ 2 Aberrations for cylindrical surfaces: generic picture
The number of aberrations for generic optical surfaces is high. The wave aberration $\Phi$, which describes the optical path difference between the actual wavefront and the intended ideal one, is a function of four parameters [12], e.g., the coordinates {$x$, $y$, $x^{\prime}$, $y^{\prime}$} of the cross-points $P_{0}(x,y,0)$ and $P^{\prime}(x^{\prime},y^{\prime},z^{\prime})$ for a paraxial ray with the object plane and the exit pupil plane, respectively. These two planes are orthogonal to the $Z$ axis (in the case of rotatoinally symmetric OS it is the symmetry axis, while in the case of discussed translationally invariant OS it can be any axis that lies in the symmetry plane and is orthogonal to the vector of translation). Expansion of $\Phi$ into the four argument power series gives 35 terms in the fourth-order part $\Phi^{(4)}$, which derivatives determine the number of the third-order aberrations (see Eqs. (8) and (9) below). In general, there are 20 different third-order power products for the four variables. Number of aberrations can be reduced for OS with definite symmetries, in particular, single-plane or double-plane symmetry. In the paraxial region a generic anamorphic OS (double-plane symmetric one) can be approximated by two independent rotationally symmetric OS, each associated with one symmetry plane [1]. This helps to derive 16 monochromatic primary aberration coefficients for anamorphic systems made from cylindrical surfaces [2].
The symmetry of a cylindrical surface, namely, single-plane symmetry plus translational invariance, allows for a simplified generic consideration and explicit analytical description presented in the following. Consider a system of cylindrical surfaces, which is symmetric under reflection in the $YZ$ plane and translationally invariant along the $Y$ axis. The translational invariance leaves three independent variables, e.g., $x$, $x^{\prime}$ and the difference $y-y^{\prime}$. We may construct from them the following six invariants of the second order
@@
The discussed plane symmetry leaves only four invariants of
@@
which completely determine the ray position. Since the variables enter only in the quadratic combinations of (2), it follows that $\Phi^{(4)}$ must be of the form
$\Phi^{(4)}=a_{1}A^{2}+\frac{a_{2}}{4}B^{2}+\frac{a_{3}}{4}C^{2}+\frac{a_{4}}{2}D^{2}$
$+\frac{a_{5}}{2}AB+\frac{a_{6}}{2}AC+a_{7}AD+\frac{a_{8}}{2}BC+a_{9}BD+a_{10}CD,$ (3)
where $a_{i}$ are the constants, and the numerical factors of $1/2$ and $1/4$ are introduced for the further convenience.
@@
where the refractive index of the material behind the exit pupil is considered equal to one for brevity, $O^{\prime}P_{1}^{*}=R^{\prime}$ is the radius of this sphere and $D_{1}$ is the distance between the exit pupil plane and the image plane. Following the consideration of Ref. [13], we take into account the fact that for the case of small aberrations in the paraxial region the directions of normal to the Gaussian reference sphere and normal to the real wave surface for the respective actual ray $P^{\prime}P_{1}$ are about the same. Then equation for the real wave surface can approximately be written as $f(x^{\prime},y^{\prime},z^{\prime})=0$ with
$f=(x^{\prime}-x_{1}^{*})^{2}+(y^{\prime}-y_{1}^{*})^{2}+(z^{\prime}-D_{1})^{2}-(s^{\prime}+\Phi)^{2},$ (5)
where $O^{\prime}P_{1}=s^{\prime}$ is the distance from $O^{\prime}$ to the actual image point $P_{1}$, and $\Phi\ll R^{\prime}\approx s^{\prime}$ with the unity refractive index for short. Equation of the normal to the real wave in the point with the coordinates $x^{\prime}$, $y^{\prime}$ and $z^{\prime}$ can be expressed as
@@
Substituting the coordinates of $P_{1}(x_{1},y_{1},D_{1})$ into (7), we get the following expressions for the components of transverse aberration
@@
$\frac{\delta_{1x}}{D_{1}}$ $=$ $(a_{2}B+a_{5}A+a_{8}C+2a_{9}D)x^{\prime}$ (10)
$+$ $(a_{4}D+a_{7}A+a_{9}B+a_{10}C)x,$
@@
$\delta_{1x}$ $=$ $(\tilde{a}_{2}x^{\prime 2}+\tilde{a}_{8}y^{\prime 2})x^{\prime}+(3\tilde{a}_{9}x^{\prime 2}+\tilde{a}_{10}y^{\prime 2})x$ (12)
$+$ $(\tilde{a}_{4}+\tilde{a}_{5})x^{2}x^{\prime}\quad+\tilde{a}_{7}x^{3},$
@@
where in the right-hand sides the first, second and third terms account for the three-component generalized spherical aberration that we call elliptic aberration for short ($e$), two-component modified coma ($c$) and astigmatism with the field curvature ($a$), respectively, while the last term in (12) represents the distortion ($d$). These aberrations are discussed one by one in the following. In total they contain 8 degrees of freedom since $\tilde{a}_{4}$ and $\tilde{a}_{5}$ parameters enter the same term. In case of $\tilde{a}_{2}=\tilde{a}_{3}=\tilde{a}_{8}$ and $\tilde{a}_{9}=\tilde{a}_{10}$ we end up with the standard set of five aberrations for the spherically symmetric OS.
@@ 2.1 Elliptic aberrations
The elliptic aberrations generalize the spherical one and depend on the three parameters $\tilde{a}_{2}$, $\tilde{a}_{3}$ and $\tilde{a}_{8}$. In the polar coordinates
@@
one can split Eqs. (16)–(17) into the two sets of parametric equations for the two ellipses:
@@
In the case of $\tilde{a}_{2}=\tilde{a}_{3}$ these two ellipses degenerate to a superposition of the two parametric circles (with the opposite circle directions), which in turn degenerate to one circle in the case of $\tilde{a}_{2}=\tilde{a}_{3}=\tilde{a}_{8}$ as shown in Fig. 1. This circle corresponds to a usual spherical aberration.
Figure 1: Elliptic aberrations in the exit pupil plane $x^{\prime}y^{\prime}$ for $\rho^{\prime}=\tilde{a}_{2}=1$, $\tilde{a}_{3}=0.5$ (left) or $\tilde{a}_{3}=1$ (right), and the denoted values of $\tilde{a}_{8}$: 1/3, 2/3, 1 and 4/3 (in unspecified units).
@@ 2.2 Modified coma
In the same coordinates defined by Eqs. (14) and (15) the terms representing modified coma can be written as
@@ 2.3 Astigmatism and field curvature
The above equations describe the ellipse (similarly to the case of spherically symmetric systems) and can be rewritten as
@@ 2.4 Distortion
It does not depend on the coordinates $x^{\prime}$ and $y^{\prime}$ of the point, in which the ray crosses the exit pupil. Similarly to the case of spherically symmetric OS, departure of the image point from the ideal image one due to the distortion is proportional to $x^{3}$.
@@ 3.1 Application of Fermat’s Principle
The homogeneous medium before the surface along the $Z$ axis has refractive index $n_{0}$ and the one after the surface has index $n_{1}$. We remark that (38) corresponds to the $XZ$ section of a generalized surface of rotation discussed in Ref. [14], where $\kappa$ is the conic constant and the $b$ term explicitly includes the type of aspheric term required for corrector plates. In particular, $\alpha=0$ for parabolic cylinder, $\alpha=R^{-3}$ for cylinder of revolution and $\alpha=(1+\kappa)R^{-3}$ with $\kappa<-1$ for hyperbolic cylinder.
@@
Figures 2 and 3 show, respectively, the $XZ$ and $YZ$ projections of the cylindrical surface of (38), whose generatrix coincides with the $Y$ axis and the symmetry axis of the aperture stop. The optical path length between the object point $P_{0}$ and the image point $P_{1}$ for the ray crossing the surface in $P^{\prime}$ can be written as
$n_{0}P_{0}P^{\prime}+n_{1}P^{\prime}P_{1}=n_{0}\sqrt{(x^{\prime}-x_{0})^{2}+y^{\prime 2}+(z^{\prime}-z_{0})^{2}}+$
$n_{1}\sqrt{(x^{\prime}-x_{1})^{2}+(y^{\prime}-y_{1})^{2}+(z^{\prime}-z_{1})^{2}},$ (39)
@@
Since normal to the surface in (38) in the points along the $Y$ axis is parallel to the $Z$ axis, the Snell’s law of refraction for the chief ray can be written as
@@
$P_{0}P^{\prime}$ $=$ $s_{0}\left[1+\frac{2x^{\prime}\sin\theta_{0}}{s_{0}}+\frac{x^{\prime 2}}{s_{0}}\left(\frac{1}{s_{0}}+\frac{\cos\theta_{0}}{R}\right)\right.$ (45)
$+$ $\left.\frac{x^{\prime 4}}{4s_{0}^{2}}\left(\frac{1}{R^{2}}+\alpha s_{0}\cos\theta_{0}\right)+\frac{y^{\prime 2}}{s_{0}^{2}}\right]^{1/2}.$
@@
$P_{0}P^{\prime}=s_{0}+x^{\prime}\sin\theta_{0}+\frac{\Theta_{0}^{x}}{2}\left(\frac{x^{\prime 2}}{s_{0}}-\frac{x^{\prime 3}}{s_{0}^{2}}\sin\theta_{0}\right)$
$+\frac{1}{2}\left(\frac{y^{\prime 2}}{s_{0}}-\frac{x^{\prime}y^{\prime 2}}{s_{0}^{2}}\sin\theta_{0}\right)-\frac{1}{8s_{0}^{3}}\left(x^{\prime 4}\Theta_{0}^{xx}-2x^{\prime 2}y^{\prime 2}\Theta_{0}^{xy}+y^{\prime 4}\right),$ (46)
@@
$\Theta_{0}^{x}=\cos\theta_{0}\left(\cos\theta_{0}+\frac{s_{0}}{R}\right),\ \Theta_{0}^{xy}=2-3\cos^{2}\theta_{0}-\frac{s_{0}}{R}\cos\theta_{0},$
$\Theta_{0}^{xx}=5\sin^{4}\theta_{0}-6\sin^{2}\theta_{0}\Theta_{0}+\Theta_{0}^{2}-\frac{s_{0}^{2}}{R^{2}}-\alpha s_{0}^{3}\cos\theta_{0},$
$\Theta_{0}=1+\frac{s_{0}}{R}\cos\theta_{0}.$ (47)
Using (44) and relation $\sin\eta\approx\tan\eta$ for small paraxial $\eta$, we express $y_{1}$ as
@@
where we ignored $O^{\prime}P^{\prime}$ with respect to $P_{0}P^{\prime}$. Using (41), we get
@@
$P^{\prime}P_{1}$ $=$ $s_{1}\left[1-\frac{2x^{\prime}\sin\theta_{1}}{s_{1}}+\frac{x^{\prime 2}}{s_{1}}\left(\frac{1}{s_{1}}-\frac{\cos\theta_{1}}{R}\right)\right.$ (51)
$+$ $\left.\frac{x^{\prime 4}}{4s_{1}^{2}}\left(\frac{1}{R^{2}}-\alpha s_{1}\cos\theta_{1}\right)+\frac{y^{\prime 2}}{s_{1}^{2}}\beta^{2}\right]^{1/2}$
$=$ $s_{1}-x^{\prime}\sin\theta_{1}+\frac{\Theta_{1}^{x}}{2}\left(\frac{x^{\prime 2}}{s_{1}}+\frac{x^{\prime 3}}{s_{1}^{2}}\sin\theta_{1}\right)$
$+$ $\frac{\beta^{2}}{2}\left(\frac{y^{\prime 2}}{s_{1}}+\frac{x^{\prime}y^{\prime 2}}{s_{1}^{2}}\sin\theta_{1}\right)$
@@
$\Theta_{1}^{x}=\cos\theta_{1}\left(\cos\theta_{1}-\frac{s_{1}}{R}\right),\ \Theta_{1}^{xy}=2-3\cos^{2}\theta_{1}+\frac{s_{1}}{R}\cos\theta_{1},$
$\Theta_{1}^{xx}=5\sin^{4}\theta_{1}-6\sin^{2}\theta_{1}\Theta_{1}+\Theta_{1}^{2}-\frac{s_{1}^{2}}{R^{2}}+\alpha s_{1}^{3}\cos\theta_{1},$
$\Theta_{1}=1-\frac{s_{1}}{R}\cos\theta_{1}.$ (52)
Now the optical path length in (3.1) can be rewritten as
$(n_{0}s_{0}+n_{1}s_{1})+x^{\prime}(n_{0}\sin\theta_{0}-n_{1}\sin\theta_{1})$
$+\frac{x^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}\Theta_{0}^{x}+\frac{n_{1}}{s_{1}}\Theta_{1}^{x}\right)-\frac{x^{\prime 3}}{2}\left(n_{0}\frac{\sin\theta_{0}}{s_{0}^{2}}\Theta_{0}^{x}-n_{1}\frac{\sin\theta_{1}}{s_{1}^{2}}\Theta_{1}^{x}\right)$
$+\frac{y^{\prime 2}}{2}\left(\frac{n_{0}}{s_{0}}+\frac{n_{1}\beta^{2}}{s_{1}}\right)-\frac{x^{\prime}y^{\prime 2}}{2}\left(\frac{n_{0}\sin\theta_{0}}{s_{0}^{2}}-\frac{n_{1}\beta^{2}\sin\theta_{1}}{s_{1}^{2}}\right)$
$-\frac{x^{\prime 4}}{8}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xx}+\frac{n_{1}}{s_{1}^{3}}\Theta_{1}^{xx}\right)+\frac{x^{\prime 2}y^{\prime 2}}{4}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xy}+\frac{n_{1}\beta^{2}}{s_{1}^{3}}\Theta_{1}^{xy}\right)$
$-\frac{y^{\prime 4}}{8}\left(\frac{n_{0}}{s_{0}^{3}}+\frac{n_{1}\beta^{4}}{s_{1}^{3}}\right).$ (53)
For the chief ray optical path length, using (42), at the same order of precision we have
$n_{0}\sqrt{x_{0}^{2}+z_{0}^{2}+y^{\prime 2}}+n_{1}\sqrt{x_{1}^{2}+z_{1}^{2}+(y^{\prime}-y_{1})^{2}}$
$=n_{0}s_{0}\sqrt{1+\frac{y^{\prime 2}}{s_{0}^{2}}}+n_{1}s_{1}\sqrt{1+\frac{(y^{\prime}-y_{1})^{2}}{s_{1}^{2}}}$
$\approx(n_{0}s_{0}+n_{1}s_{1})+\frac{n_{0}y^{\prime 2}}{2s_{0}}-\frac{n_{0}y^{\prime 4}}{8s_{0}^{3}}+\frac{n_{1}\beta^{2}y^{\prime 2}}{2s_{1}}-\frac{n_{1}\beta^{4}y^{\prime 4}}{8s_{1}^{3}}.$
By subtracting (3.1) from (3.1) we get the wave aberration
$\Phi^{(4)}=\alpha_{d}\,x^{\prime}+\alpha_{a1}\frac{x^{\prime 2}}{2}+\alpha_{a2}\frac{y^{\prime 2}}{2}+\alpha_{c1}\frac{x^{\prime 3}}{3}+\alpha_{c2}\,x^{\prime}y^{\prime 2}$
$+\leavevmode\nobreak\ \alpha_{e1}\frac{x^{\prime 4}}{4}+\alpha_{e2}\frac{y^{\prime 4}}{4}+\alpha_{e3}\frac{x^{\prime 2}y^{\prime 2}}{2},$ (55)
where the terms with the coefficients $\alpha_{a2}$ and $\alpha_{e2}$ vanish due to the approximations used, while the rest coefficients read
$\alpha_{e1}$ $=$ $-\frac{1}{2}\left(\frac{n_{0}}{s_{0}^{3}}\Theta_{0}^{xx}+\frac{n_{1}}{s_{1}^{3}}\Theta_{1}^{xx}\right),$ (56)